OpenStudy (anonymous):

solve linear and literal equations. solve for F c=(5/9)(F-32)

OpenStudy (anonymous):

1) multiply both sides by \(\frac{9}{5}\) 2) add 32

OpenStudy (anonymous):

\[C=\frac{5}{9}(F-32)\] \[\frac{9}{5}C=F-32\] \[\frac{9}{5}C+32=F\]

OpenStudy (anonymous):

thank you soooo much!

OpenStudy (anonymous):

yw, hope it is clear